An object of mass m = 50 g is released from rest from a height h = 150 m above the ground. Calculate the kinetic energy and the potential energy of the object at the height H = h/5.
1
Expert's answer
2020-03-19T11:30:48-0400

The total energy of an object at the initial point


"E=KE+PE=0+mgh""=0.05\times 9.8\times 150=73.5\:\rm J"

At the height "H=h/5=30\:\rm m," the potential energy of an object


"PE=mgH=0.05\times 9.8\times 30=14.7\:\rm J"

The kinetic energy


"KE=E-PE=73.5-14.7=58.8\:\rm J"